### v1

# Dark Dimension and the Effective Field Theory limit

Abstract
Abstract
In [1] we pointed out that in the Dark Dimension scenario [2] theoretical issues arise when the prediction for the vacuum energy $\rho$, that is obtained from swampland conjectures in string theory, is confronted with the corresponding result for $\rho$ in the field theory limit. One of the problems arises from the field theory side and is due to previous results for $\rho$ that, according to our analysis, turn out to be incorrect, although widely spread. Our work was challenged in [3]. Here we show why in our opinion the claims in [3] are inconsistent, actually based on arguments deprived of physical meaning, providing further support to our findings. We conclude by discussing ideas on the physical mechanism that could trigger the suppression of large UV contributions to $\rho$.
In a recent paper [1], we analysed aspects of the dark dimension (DD) proposal [2], according to which above the neutrino scale our universe is described by a $(4+1)$D theory with one compact dimension of micrometer size, and pointed out possible theoretical issues. Our work was challenged in [3]. Here we show why in our opinion the authors of [3] (in their attempt to confute our results) bring confusion on the whole subject, and use incorrect arguments that lead them to physically meaningless conclusions. In view of the great wealth of theoretical and phenomenological work triggered by the DD scenario, we believe it worth trying to bring clarity on the subject and continuing its investigation.
To this end, we begin by recalling the main points [^1] For convenience of the reader less familiar with the DD proposal, in Appendix A we provide a more detailed summary of the different aspects and steps that lead to this scenario. of the DD scenario [2] together with the main results of our work [1]. In a nutshell, the DD proposal combines swampland conjectures with observational data and phenomenological bounds, and is based on the assumption that our universe lies in a unique (asymptotic) corner of the quantum gravity landscape, where only two kinds of light tower states are expected: towers of string excitations or towers of KK states. Indicating with $\mu_{tow}$ the tower scale, and with $\Lambda_{cc}$ the measured vacuum energy (cosmological constant times $M_{P}^{2}$, with $M_{P}$ the Planck mass), the application of the distance conjecture to (A)dS vacua relates the two as
$\mu_{tow}\sim\left|\frac{\Lambda_{cc}}{M_{P}^{4}}\right|^{\alpha}M_{P}\,.$ (1)
Furthermore, combining indications from one-loop string calculations with the Higuchi bound [4], the proponents of the DD scenario restrict the parameter $\alpha$ to the range
$\frac{1}{4}\leq\alpha\leq\frac{1}{2}\,.$ (2)
Then, considering bounds set by the non-observation of deviations from Newton’s force law, they infer that $\mu_{tow}$ is close to the energy scale associated to the cosmological constant (that coincides with the neutrino scale)
$\mu_{tow}\sim\Lambda_{cc}^{1/4}\sim\,\,{\rm meV}\,.$ (3)
They also observe that, for (1) to be consistent with (3), it is necessary to take $\alpha$ at the lower bound of (2), $\alpha=1/4$.
The authors then make the following crucial observation. Let us quote from [2]: “since we can describe physics above the neutrino scale with effective field theory” (EFT), the string scenario is “ruled out experimentally”. They then conclude that the light tower must be a KK tower, $\mu_{tow}=m_{{}_{\rm KK}}$, so that (1) becomes
$\Lambda_{cc}\sim m_{{}_{\rm KK}}^{4}\,.$ (4)
Finally, to determine the number of extra compact dimensions, they compare bounds on $m_{{}_{\rm KK}}$ that come from calculations of the heating of neutron stars due to the surrounding cloud of trapped KK gravitons [5, 6] with the value of $m_{{}_{\rm KK}}$ in (4). This allows them to conclude that there is only one extra dimension of size $\sim\mu{\rm m}$.
The authors of [3] start their rebuttal of our work [1] claiming that we question the basic (swampland) relation (4) of the DD proposal. Before going into more details, we stress that in [1] we do not question (4), nor was this our goal. What we did is different. We now briefly recall and clarify the scope and the main results of [1].
As already said, according to the proponents of the DD scenario, above the $m_{{}_{\rm KK}}$ scale our universe is described by a $(4+1)$D EFT with one compact dimension of micrometer size. This proposal has triggered a large amount of work, where $(4+1)$D EFT models that could realize the DD scenario are proposed, and their phenomenological consequences studied (see for instance [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]). In [1], our goal was to investigate on the matching between the string theory framework in which the basic relation (4) of the DD proposal is obtained and its low energy EFT limit.
To this end, we considered a supersymmetric (4+1)D theory with Scherk-Schwarz SUSY breaking (and also a non-supersymmetric theory) with compact dimension in the shape of a circle [^2] In the DD scenario the Standard Model (SM) lives on a 3-brane, and it would be then more appropriate to consider an orbifold. For our scope, namely the calculation of the one-loop contribution of bulk fields to the vacuum energy, it is sufficient to consider the simplest case of a circle. of radius $R$ and calculated the one-loop vacuum energy $\rho^{1l}$. According to existing literature (see for instance [21, 20, 19, 18], and for the analogous calculation of the effective potential [22, 23]), in $(4+n)$D theories with $n$ compact dimensions $\rho^{1l}$ should be finite and proportional to $m_{{}_{\rm KK}}^{4}$,
$\rho^{1l}\sim m_{{}_{\rm KK}}^{4}\,,$ (5)
with no need for fine-tuning. In [1] we showed that the right hand side of (5) also contains UV-sensitive terms (see the first line in the right hand side of Eq. (Dark Dimension and the Effective Field Theory limit) below), that were missed in previous literature. Therefore, the matching between the prediction (4) for the vacuum energy $\Lambda_{cc}$, that comes from swampland conjectures in string theory, and the corresponding result in the low energy field theory limit is not as straightforward as it was thought on the basis of the result (5). In this respect, it is worth to stress again (see the paragraph below Eq. (3)) that such a matching is crucial for the DD scenario.
In [1] we show why it is commonly thought that $\rho^{1l}$ is finite and proportional to $m_{{}_{\rm KK}}^{4}$: this result comes from a mistreatment of the loop momentum asymptotics. Performing the calculation properly, we found that (in addition to the finite term $\sim m^{4}_{{}_{\rm KK}}$ and to divergences that cancel out in SUSY theories) the one-loop contribution to the vacuum energy $\rho^{1l}$ contains “UV-sensitive” terms proportional to powers of the boundary charges. These divergences do not disappear, not even in a SUSY theory. This is due to the fact that, to trigger the Scherk-Schwarz mechanism, the boundary charges $q_{b}$ and $q_{f}$ of the boson and fermion superpartners need to be different. Below we show how the authors of [3] misuse and distort our findings, and how this leads them to meaningless conclusions.
Our results for $\rho^{1l}$ are in Eqs. (25) and (26) of [1]. For ease of presentation (and with no loss of generality), from now on we concentrate on the SUSY theory case only (unless expressly indicated otherwise), and write the contribution to $\rho^{1l}$ from a single couple of superpartners for the massless case (i.e. taking $m=0$ in (25) and (26) of [1])
$\rho^{1l}$ $=\frac{(q^{2}_{b}-q^{2}_{f})\,R\,m_{{}_{\rm KK}}^{2}}{60\pi^{2}}\Lambda^{3}-\frac{(q_{b}^{4}-q_{f}^{4})\,R^{\frac{1}{3}}\,m_{{}_{\rm KK}}^{\frac{10}{3}}}{280\pi^{2}}\Lambda$
+ $+\frac{\,3\left[\text{Li}_{5}(e^{2\pi iq_{b}})-\text{Li}_{5}(e^{2\pi iq_{f}})+h.c.\right]\,}{128\pi^{6}}\,m_{{}_{\rm KK}}^{4}\,,$ (6)
where $q_{b}$ and $q_{f}$ are the boson and fermion boundary charges respectively, and $\text{Li}_{5}(x)$ is the polylogarithm function $\text{Li}_{n}(x)$ for $n=5$. $\Lambda$ is the physical UV cutoff of the $(4+1)$D theory defined by the action $\mathcal{S}^{(4+1)}=\mathcal{S}_{\rm grav}^{(4+1)}+\mathcal{S}_{\rm{matter}}^{(4+1)}$, with $(4+1)$D metric $\hat{g}_{{}_{MN}}$ parametrized in terms of the graviphoton $A_{\mu}$ and the radion field $\phi$ (see Appendix B, and equations (6)-(12) in [1]). The KK scale $m_{{}_{\rm KK}}$ is given in terms of the radion field vev (here indicated with $\phi$)
$m_{{}_{\rm KK}}=e^{\sqrt{\frac{3}{2}}\frac{\phi}{M_{P}}}R^{-1}\equiv R_{\phi}^{-1}\,.$ (7)
To be phenomenologically viable, the $(4+1)$D model that implements the DD scenario has to include the SM on a 3-brane. An important consequence is that the physical cutoff $\Lambda$ of the $(4+1)$D theory is an upper bound for the SM cutoff $\Lambda_{\rm SM}$ (the derivation of Eq. (8) below is in Appendix B)
$\Lambda_{\rm SM}\leq\Lambda\,.$ (8)
This result is important for our following analysis. Let us also note that $m_{{}_{\rm KK}}$ is the cutoff of the “reduced 4D theory”, which besides brane fields contains the zero modes of bulk fields. In Fig. 1 we provide a comprehensive picture of the general framework in which the DD scenario is framed. Finally, it is worth to observe that, although both the Standard Model and this “reduced theory” are $4$-dimensional EFTs, they are certainly not the same theory, and their cutoffs ($\Lambda_{\rm SM}$ and $m_{{}_{\rm KK}}$ respectively) should not be confused.
In Eq. (Dark Dimension and the Effective Field Theory limit), the divergences (i.e. UV-sensitive terms) are in the first line, the finite term in the second one. Limiting ourselves to consider only the most divergent term of $\rho^{1l}$ (and doing the same for the non-SUSY theory, whose result for $\rho^{1l}$ is not reported above but can be found in our work [1]), we have
$\rho^{1l}\sim m_{{}_{\rm KK}}^{2}R\Lambda^{3}\qquad{\rm and}\qquad\rho^{1l}\sim m_{{}_{\rm KK}}^{2/3}R^{5/3}\Lambda^{5}\,,$ (9)
for the SUSY and the non-SUSY case respectively. These latter equations are Eqs. (28) and (29) of [1]. As stressed above, one of the main goals of our work [1] is to point out that $\rho^{1l}$ contains previously unnoticed divergences (that do not disappear even in a SUSY theory) and that, before comparing $\rho^{1l}$ with the measured value of the vacuum energy $\Lambda_{cc}$, a fine-tuning is needed.
Figure 1: Pictorial representation of the energy scales and theories involved in the Dark Dimension scenario. The grey region indicates a possible UV-completion of the $(4+1)$D theory between its (rescaled, see text) physical cutoff $\Lambda_{\phi}$ and the string scale $M_{s}$. A particular case could be that $\Lambda_{\phi}$ directly coincides with $M_{s}$. Below $\Lambda_{\phi}$ (that coincides with the Standard Model physical cutoff $\Lambda_{\rm SM}$, see text) the $(4+1)$D theory of the DD scenario (blue region), that contains bulk fields and SM (with possibly other) brane fields, takes place. Below $m_{{}_{\rm KK}}$, the theory is well described by a $4D$ EFT (green region) containing the 00-modes of the bulk fields together with brane fields. The components $p$ and $p_{5}$ of the $(4+1)$D momentum $\hat{p}\equiv(p,p_{{}_{5}})$, with $p\equiv(p_{1},p_{2},p_{3},p_{4})$ and $p_{5}$ the fifth component, are reported. The fully colored blue and green rectangles indicate that $p$ takes continuous values (non-compact dimensions), while the blue lines indicate that $p_{5}$ takes discrete values starting at $m_{{}_{\rm KK}}$ (compact dimension).
For the purposes of the present analysis, it is of fundamental importance to stress that the authors of  [3] write our Eqs. (9) as (see Eqs. (4) of [3])
$\Lambda_{cc}\sim m_{{}_{\rm KK}}^{2}R\Lambda^{3}\qquad{\rm and}\qquad\Lambda_{cc}\sim m_{{}_{\rm KK}}^{2/3}R^{5/3}\Lambda^{5}\,,$ (10)
thus operating in (9) the replacement
$\rho^{1l}\to\Lambda_{cc}\,.$ (11)
So doing, they take the result $\rho^{1l}$ of our one-loop calculation in [1] to directly coincide (without any fine-tuning) with the measured value of the vacuum energy.
Although this is opposite to what we did (we repeat ourselves, in [1] our goal was to point out the presence of unexpected UV-sensitive terms), the replacement (11) is not in itself a problem. It is a theoretically legitimate possibility that the cutoff $\Lambda$ is a physical scale that appears in physical quantities, with no need for fine-tuning. However, the consequences of such a choice have to be fully taken into account. Anticipating on the results, we will show that the replacement $\rho^{1l}\to\Lambda_{cc}$ has crucial physical consequences that the authors of [3] failed to recognize. It is precisely this replacement that leads them to meaningless conclusions. Let us note that, for reasons that will be soon apparent, the right hand side of (10)1 should contain all the terms that appear in the right hand side of (Dark Dimension and the Effective Field Theory limit), not only the first one. However, since this has no impact on the present discussion, we can equivalently refer to (10)1 or more completely to (Dark Dimension and the Effective Field Theory limit).
It is now crucial to realise that, once $\rho^{1l}$ is chosen to coincide with the measured vacuum energy $\Lambda_{cc}$, Eqs. (10) completely determine $\Lambda$. As we will see, the authors of [3] base their analysis on Eqs. (10) and, failing to realise that they do not have this freedom, freely play with the cutoff $\Lambda$. Below we determine the value of $\Lambda$ imposed by their choice (11), and investigate on the possibility that the resulting $\Lambda$ can really play the role of the theory UV cutoff. For this to be true, $\Lambda$ cannot violate well-tested phenomenology, in particular it should not enter in conflict with the Standard Model.
Let us determine now the value of $\Lambda$ imposed by (11). First of all, we recall that from (4) we have $m_{{}_{\rm KK}}\sim\Lambda_{cc}^{1/4}$. A simple inspection of  (10)1 (or (Dark Dimension and the Effective Field Theory limit) with the replacement (11)) shows that this implies [^3] The same is true for (10)2.
$R\Lambda^{3}\sim m^{2}_{{}_{\rm KK}}\sim({\rm meV})^{2}\,.$ (12)
The above relation is the unavoidable consequence of  (11). Starting from (12), it is now easy to show that the whole construction put forward in [3] is flawed. In fact, using the relation ($\hat{M}_{P}$ is the five-dimensional Planck scale)
$\hat{M}_{P}^{3}=(2\pi R)^{-1}M_{P}^{2}\,,$ (13)
Eq. (12) is straightforwardly rewritten as
$\Lambda^{3}\sim\left(\frac{m^{2}_{{}_{\rm KK}}}{M_{P}^{2}}\right)\hat{M}_{P}^{3},$ (14)
from which we obtain (remember that $m_{{}_{\rm KK}}\sim$ meV)
$\Lambda\,\sim\,10^{-21}\hat{M}_{P}\,\leq\,10^{-2}\,\text{GeV},$ (15)
where the last inequality stems from $\hat{M}_{P}\leq M_{P}$. Eq.(15) seems to set a too low upper bound on the cutoff $\Lambda$ of the $(4+1)$D theory. We will see in a moment that this is really the case.
The problem comes from the relation between $\Lambda$ and the $4$D cutoff $\Lambda_{\rm SM}$ of the Standard Model, that, as already said, lives on a 3-brane. In Appendix B (to which we refer for the calculation) we show that $\Lambda_{\rm SM}$ is given by
$\Lambda_{\rm SM}=\Lambda_{\phi}\equiv e^{\frac{1}{\sqrt{6}}\frac{\phi}{M_{P}}}\Lambda,$ (16)
where $\phi$ is the $vev$ of the radion field (see comments above Eq. (7) and Appendix B) and $\Lambda_{\phi}$ is the $(4+1)$D cutoff for the momenta rescaled by the factor $e^{\frac{1}{\sqrt{6}}\frac{\phi}{M_{P}}}$, that are nothing but the momenta from the $4$D observer perspective (see Eq. (47) in Appendix B). Since the asymptotic scenario of the DD proposal is realized for negative values of $\phi$ (see Appendix A), Eqs. (15) and (16) imply
$\Lambda_{\rm SM}<10^{-2}\,{\rm GeV}\,,$ (17)
an unacceptably low upper bound on $\Lambda_{\rm SM}$.
We can go further and determine the value that $\Lambda_{\rm SM}$ should have if the replacement $\rho^{1l}\to\Lambda_{cc}$ operated by the authors of [3] could be accepted. In fact, observing that the species scale $\Lambda_{\rm sp}$[24, 25] is given by
$\Lambda_{\rm sp}\sim m_{{}_{\rm KK}}^{1/3}M_{P}^{2/3}\sim e^{\frac{1}{\sqrt{6}}\frac{\phi}{M_{P}}}\hat{M}_{P}$ (18)
(see for instance [26] and the Appendix of [1]), comparing (16) with (18) and using (15) we immediately have
$\Lambda_{\rm SM}\sim\frac{\Lambda}{\hat{M}_{P}}\Lambda_{\rm sp}\sim 10^{-21}\Lambda_{\rm sp}.$ (19)
Finally, since from (18) we have $\Lambda_{\rm sp}\sim 10^{9}$ GeV,
$\Lambda_{\phi}=\Lambda_{\rm SM}\sim 10^{-12}{\rm GeV}=10^{-3}{\rm eV}\,,$ (20)
a too low value, as already known from (17).
Eq. (20) definitely shows that what the authors of [3] do when they present our Eqs. (9) as Eqs. (10) (that are nothing but (9) with the replacement $\rho^{1l}\to\Lambda_{cc}$), is physically not allowed. They operate an inconsistent replacement, and Eqs. (10) are deprived of any physical meaning. The point is that the arguments they use in their attempt to confute our work completely depend on this replacement, actually cannot even be formulated without it. Their arguments are thus unfounded and inconsistent.
There is a simpler way to obtain (20) for $\Lambda_{\rm SM}$. From (7) and (16) it is easily seen that $R\Lambda^{3}=R_{\phi}\Lambda_{\phi}^{3}$, so that Eqs. (10) can be written as
$\Lambda_{cc}\sim m_{{}_{\rm KK}}^{2}R_{\phi}\Lambda_{\phi}^{3}\qquad{\rm and}\qquad\Lambda_{cc}\sim m_{{}_{\rm KK}}^{2/3}R_{\phi}^{5/3}\Lambda_{\phi}^{5}\,.$ (21)
Therefore, inserting  $R_{\phi}=m_{{}_{\rm KK}}^{-1}$ in (21), we immediately find that $\Lambda_{cc}\sim m_{{}_{\rm KK}}^{4}$ is obtained only if
$\Lambda_{\phi}=\Lambda_{\rm SM}\sim m_{{}_{\rm KK}}\sim 10^{-3}\,{\rm eV}\,,$ (22)
that is nothing but (20). Eq. (22) (and equivalently Eq. (20)) shows a further drawback of the (illegitimate) replacement $\rho^{1l}\to\Lambda_{cc}$. From this equation in fact we see that the latter forces the $(4+1)$D cutoff $\Lambda_{\phi}$ to be squeezed upon $m_{{}_{\rm KK}}$. This leaves no space to the $(4+1)$D EFT needed for the DD scenario, and one is only left with a $4$D theory valid below $m_{{}_{\rm KK}}$ ($\sim$ meV), above which the theory directly merges to its UV completion (see Fig. 1, where the choice $\Lambda_{\phi}=m_{{}_{\rm KK}}$ makes the blue region disappear). Therefore, if the replacement $\rho^{1l}\to\Lambda_{cc}$ operated in [3] could be accepted, this would close the door to the DD scenario. Certainly, this was not the original intention of the authors in [3].
The above analysis shows that Eqs. (10) are physically inconsistent, and obviously no use can be made of them. In particular, they cannot be used to confute our results. Unfortunately, this is precisely what the authors of [3] do, thus developing arguments and claims that are clearly deprived of any physical meaning.
The present work could end here. Nevertheless, in the hope that this can bring further clarity and help the reader gain more insight in the whole subject, we now provide a detailed analysis of their arguments, that (as already said) are based on the inconsistent Eqs. (10). To follow their presentation closely, below we consider both the SUSY and the non-SUSY cases, i.e. both their Eqs. (10)1 and (10)2.
One of the authors’ first points is to compare the (A)dS distance conjecture (1), that for the reader’s convenience we rewrite here as ($\mu_{tow}$ in (1) is clearly $m_{{}_{\rm KK}}$)
$m_{{}_{\rm KK}}\sim M_{P}^{1-4\alpha}\,\Lambda_{cc}^{\alpha}\,,$ (23)
with Eqs. (10), that again for convenience we write as
$m_{{}_{\rm KK}}\sim(R\Lambda^{3})^{-1/2}\Lambda_{cc}^{1/2}\,\,\,;\,\,\,m_{{}_{\rm KK}}\sim(R\Lambda^{3})^{-5/2}\Lambda_{cc}^{3/2}\,.$ (24)
This comparison leads them to affirm that for the parameter $\alpha$ we claim the values $\alpha=1/2$ and $\alpha=3/2$, for the SUSY and non-SUSY case respectively.
Nothing could be further from us. Such, allegedly ours, claim actually results from (i) the replacement (11) that the authors of [3] operate in our equations (9), thus obtaining Eqs. (10), and (ii) the comparison of these latter equations with (23). Since we know that Eqs. (10) cannot be written (simply do not exist), we never could have made such a claim. Having clarified this point, we further observe that even if for a moment we follow the authors’ line of reasoning, it is immediate to see from (23) and  (24) that the values $\alpha=1/2$ and $\alpha=3/2$ (for the SUSY and non-SUSY case, respectively) can be derived only if it is also verified that
$R\Lambda^{3}\sim M_{P}^{2}\,.$ (25)
Closing for a moment our eyes on the phenomenological impossibility of having (20), what we have just found would mean that, for (1) and  (10) to give $\alpha=1/2$ and $\alpha=3/2$ (for the SUSY and non-SUSY case, respectively), Eqs. (12) and (25) should both be fulfilled at the same time, which is obviously an absurd requirement.
Let us move now to another of the arguments developed in [3]. The authors consider the Higuchi bound  [^4] We remind here that according to the Higuchi bound the physical mass $M$ of a spin-2 particle in dS space and the physical value of the cosmological constant satisfy [4], $M^{2}\geq\frac{2}{3}\frac{\Lambda_{cc}}{M_{p}^{2}}\,.$ (26) , that in the present case reads
$\Lambda_{cc}\leq\left(\frac{3}{2}M_{P}^{2}\right)m_{{}_{\rm KK}}^{2}\,,$ (27)
and comparing (27) with (10) they get
$R\Lambda^{3}\lesssim M_{P}^{2}\qquad{\rm and}\qquad R\Lambda^{3}\lesssim 10^{-48}M_{P}^{2}$ (28)
for the SUSY and non-SUSY cases, respectively [^5] To obtain (28)2, $\Lambda_{cc}\sim 10^{-120}M_{P}^{4}$ is used.. As we know, Eqs. (28) would not even exist without the inconsistent replacement $\rho^{1l}\to\Lambda_{cc}$ in our equations (9). Nevertheless, closing again an eye on this point, we now follow their reasoning. Concerning (28)1, they observe that this relation does not constrain $R\Lambda^{3}$ too strongly, while for (28)2 they say it imposes a too strong constraint on $\Lambda$. This latter claim deserves a comment. Using the relation (13), Eq. (28)2 can be written as
$\Lambda\lesssim 10^{-16}\,\hat{M}_{P}\,.$ (29)
At the same time, using again $R\Lambda^{3}=R_{\phi}\Lambda_{\phi}^{3}$ (see comments above Eq.(21)), Eq. (28)2 can also be written in terms of $\Lambda_{\phi}$ as
$\Lambda_{\rm SM}=\Lambda_{\phi}\lesssim 10^{5}\,m_{{}_{\rm KK}}\sim 10^{2}\,{\rm eV}.$ (30)
A simple look at (29) and (30) shows that the trouble with (28)2 is not that it implies a too low value for the cutoff $\Lambda$ (as stated in [3]), but rather that it implies a too low value for the SM cutoff $\Lambda_{\rm SM}$, as we see from (30). Yet another contradiction is found if we observe that, combining (16) with (30), at threshold (i.e. when in (30) $\Lambda_{\phi}\sim 10^{2}$ eV) we have
$\Lambda\gtrsim 10^{2}\,{\rm eV}.$ (31)
In fact, few lines after their claim that (28)2 implies a too low value for $\Lambda$, the authors of [3] make another claim, namely that the “correct choice” for $\Lambda$ is $\Lambda=m_{{}_{\rm KK}}$ (see below). As we have just seen, however, at threshold (28)2 implies (31), while $\Lambda=m_{{}_{\rm KK}}$ means $\Lambda\sim$ meV, that is five orders of magnitude below the lower bound in (31). The patent contradiction between these two claims of theirs can hardly be missed.
Successively, in [3] the authors make yet another use of their Eqs. (10). They observe that these equations still contain the radius $R$, and claim that the latter should be replaced by $m_{{}_{\rm KK}}^{-1}$, so that additional powers of $m_{{}_{\rm KK}}$ appear in these equations. Before proceeding with our analysis, we must observe that, so doing, the authors miss the crucial radion dependence in the relation between $R$ and $m_{{}_{\rm KK}}$ (see (7)).
Anyway, let us go on and follow their arguments. Invoking the UV-IR mixing occurring in string theory and/or quantum gravity, the authors of [3] say that the UV cutoff $\Lambda$ of an EFT coupled to quantum gravity should depend on the IR scale, that in this case is $m_{{}_{\rm KK}}$. They then observe that the largest possible cutoff is the species scale $\Lambda_{\rm sp}\sim m_{{}_{\rm KK}}^{1/3}M_{P}^{2/3}$ (see (18)), which in fact depends on $m_{{}_{\rm KK}}$. Now, replacing in Eqs. (10) $R$ with $m_{{}_{\rm KK}}^{-1}$ and $\Lambda$ with $\Lambda_{\rm sp}$, they finally get [^6] Actually $\Lambda_{\rm sp}$ is a cutoff for the 4D momentum $p$ and for the KK tower, (see for instance [26] and the Appendix of [1]), while $\Lambda$ is the cutoff for the $(4+1)$D momentum $\hat{p}$, that gives rise to the cutoff $\Lambda_{\phi}$ for $p$, see Eq. (47) in Appendix B. Therefore, contrary to what is done in [3], $\Lambda_{\rm sp}$ has to be identified with $\Lambda_{\phi}$ not with $\Lambda$. However, since $R\Lambda^{3}=R_{\phi}\Lambda_{\phi}^{3}$ (see comments above (21)), the two mistakes made by the authors in replacing $R$ (rather than $R_{\phi}$) with $m_{{}_{\rm KK}}^{-1}$ and $\Lambda$ (rather than $\Lambda_{\phi}$) with $\Lambda_{\rm sp}$ cancel between one another. Therefore, the correct substitutions lead to the same result (32). Moreover, we note that $\Lambda_{\rm sp}$ is the maximal value for $\Lambda_{\phi}$, and is obtained for $\Lambda=\hat{M}_{P}$, which is of course the maximal possible value for $\Lambda$.
$\Lambda_{cc}\sim m_{{}_{\rm KK}}^{2}M_{P}^{2}\qquad{\rm and}\qquad\Lambda_{cc}\sim m_{{}_{\rm KK}}^{2/3}M_{P}^{10/3}\,.$ (32)
From these results, they conclude that this first of their “UV-IR mixing motivated” attempts is not satisfactory, as (32)2 violates the Higuchi bound.
Then, still invoking the “UV-IR mixing”, they proceed by making another trial for $\Lambda$, that they eventually dub as the “correct choice”. In this second attempt they insert in Eqs. (10) $\Lambda=m_{{}_{\rm KK}}$ (and $R=m_{{}_{\rm KK}}^{-1}$, as for the previous case). A first thing worth to note in connection with this new “trial” is that a simple look to Eqs. (10) shows that, since they are fixing all the scales to $m_{{}_{\rm KK}}$, the right hand side of these equations is trivially proportional to $m_{{}_{\rm KK}}^{4}$. The authors of [3] seem to be satisfied by this result, and claim that with this “correct choice” of $\Lambda$ they are able to remove what they call the “ambiguity” in the EFT calculation and fix what they consider as the unique cutoff required by UV-IR mixing. In our opinion, this is (at least) too cheap a way to get the result “vacuum energy $\sim m_{{}_{\rm KK}}^{4}$”, and to implement the string UV-IR connection in the field theory limit. But this is not the worst. What certainly kills the possibility of taking $\Lambda=m_{{}_{\rm KK}}$ is what we have already seen. This choice is not allowed as it implies a ridiculously small cutoff for the Standard Model [^7] It is probably worth to note that if instead of the authors’ replacement $R=m_{{}_{\rm KK}}^{-1}$ one makes the correct replacement $R_{\phi}=m_{{}_{\rm KK}}^{-1}$ (see (7)), from (32) we have that $\Lambda_{cc}\sim m_{{}_{\rm KK}}^{4}$ is obtained for $\Lambda_{\phi}=m_{{}_{\rm KK}}$ rather than $\Lambda=m_{{}_{\rm KK}}$. In any case, both $\Lambda_{\phi}=m_{{}_{\rm KK}}$ and $\Lambda=m_{{}_{\rm KK}}$ are unacceptably low..
We finally comment on two additional arguments that the authors of [3] use to lend further support to their views. Making reference to the finiteness of string theory calculations, their goal is to give examples where finite results are obtained also in field theory. To this end, they consider (i) the calculation of the one-loop free energy $F^{1l}$ in finite temperature QFT, where the $T$-dependent contribution goes like $F_{T}^{1l}\sim T^{4}$ with no need for any “string regularization” (to use their words), and (ii) the calculation of the Casimir energy.
Let us begin with the finite temperature field theory case. As stressed in [27], despite their apparent similarity, there is a profound difference between the sum over the integer $n$ that appears in $F^{1l}$ and the analogous sum in the calculation of the vacuum energy $\rho^{1l}$ in KK theories. In fact, both for $\rho^{1l}$ and $F^{1l}$ we have an expression of the kind (below $d=4$ for the $(4+1)$D KK theory and $d=3$ for finite temperature field theory)
$\rho^{1l}\,\,,\,\,F^{1l}\,\sim\frac{1}{2}\,\sum_{n}\int\,d^{d}p\,\log(p^{2}+m^{2}+f_{n})\,.$ (33)
For the case of KK theories ($\rho^{1l}$) the integer $n$ is related to the fifth component of the $(4+1)$D loop momentum $\hat{p}$, that is cut at $\Lambda$. Therefore $n$ and $p$ are intertwined between one another and are to be cut accordingly [1, 27]. This in particular implies an upper bound on the integer $n$. As a result, our Eqs. (25) and (26) of [1], here reported for the massless SUSY case in Eq. (Dark Dimension and the Effective Field Theory limit), are obtained. For the finite temperature field theory case ($F^{1l}$), on the contrary, the integer $n$ and the $3D$ momentum $p$ are not intertwined. The $d^{3}p$ integral provides the trace over quantum fluctuations (and is cut at a certain momentum cutoff), while the $\sum_{n}$ gives the statistical average (mixed states), and as such has to be performed all the way up to infinity to implement ergodicity. This infinite sum is at the origin of the finiteness [^8] It is worth to note that the full result for $F^{1l}$ is finite only if the theory is supersymmetric. of the T-dependent contribution $F_{T}^{1l}$ to $F^{1l}$. In fact,
$F_{T}^{1l}$ $=\frac{T}{2}\sum_{n=-\infty}^{\infty}\int d^{3}p\,\log(p^{2}+m^{2}+f_{n})$
$-\frac{1}{2}\int d^{3}p\,\sqrt{p^{2}+m^{2}}\sim T^{4}={\rm finite}\,.$ (34)
As mentioned above and discussed at length in [1] and [27], performing in the KK case too the sum over $n$ up to infinity (while keeping $p$ finite) mistreats the loop momentum. This is what causes the (artificial) disappearance of the $q$-dependent UV-sensitive terms in (Dark Dimension and the Effective Field Theory limit), giving the impression that there is a cancellation similar to what happens in the finite temperature case.
Moving now to the example of the Casimir energy $\mathcal{E}_{C}$, we observe that the latter is nothing but the difference between the vacuum energy $\rho_{{}_{R}}$ calculated in the KK theory with one compact dimension (that for our illustrative purposes we took as a circle of radius $R$) and the corresponding $\rho_{{}_{\infty}}$ calculated in the decompactification limit (i.e.  in the theory with all non-compact dimensions): $\mathcal{E}_{C}=\rho_{{}_{R}}-\rho_{{}_{\infty}}$. Now, if we perform the one-loop calculation of $\rho^{1l}$ according to previous literature, meaning that in (33) we perform the sum over $n$ up to infinity, $\mathcal{E}_{C}^{1l}$ takes a form in all similar to (Dark Dimension and the Effective Field Theory limit), with $T^{4}$ replaced by $m_{{}_{\rm KK}}^{4}$. However, we have shown in [1] that (contrary to what is found in previous literature) the subtraction $\rho_{{}_{R}}-\rho_{{}_{\infty}}$ is not in general sufficient to ensure the disappearance of all the UV-sensitive terms. This is for instance the case of KK theories with Scherk-Schwarz SUSY breaking considered in the present work.
In summary, we have shown that the criticisms raised by the authors of [3] against our work [1] are unfounded. Their objections rely on Eqs. (10), that in turn arise from the replacement (11), namely $\rho^{1l}\to\Lambda_{cc}$, operated by the authors of [3] in our original results for $\rho^{1l}$ (here reported in Eqs. (9)). We have shown that  (11) is inconsistent and leads to arguments deprived of physical meaning. The most striking consequence of the replacement (11) is that for the physical cutoff $\Lambda_{\rm SM}$ of the Standard Model it necessarily implies $\Lambda_{\rm SM}\sim m_{{}_{\rm KK}}\sim 10^{-3}{\rm eV}$ (see Eq. (20)), which is obviously absurd. This crucial consequence of their replacement, that makes their arguments inconsistent, was overlooked by the authors of [3]. Moreover, we have shown that, even closing an eye on this severe shortcoming, the arguments presented by the authors of [3] have further flaws and lack of internal coherence.
Before ending this Letter, we would like to make a couple of additional comments on some open questions connected with the subject discussed in the present work.
Concerning the UV-IR mixing invoked in [3], there are several works where the impact that modular invariance in string theory can have on its low energy EFT limit is studied (see for instance [28, 30, 31, 29]), and it is certainly of the greatest importance to continue the investigation in these directions. However, we think that such a delicate issue cannot be trivially shortcut with the choice $\Lambda=m_{{}_{\rm KK}}$ considered in [3] and discussed above.
Finally, we would like to spend one word on the general framework in which in our opinion the DD scenario should be framed. In [1] and in the present work we considered the contribution to the vacuum energy that comes from the bulk. The reason is that in the paper that contains the DD proposal [2] as well as in the paper [3] that attempts to confute the results of our work [1] this is the only part of the vacuum energy taken into account. Naturally, there is also the contribution coming from the Standard Model, that lives on a 3-brane. Since the SM physical cutoff $\Lambda_{\rm SM}$ is at least of the TeV order, the latter cannot be ignored and viability of the DD scenario is subject to the existence of a physical mechanism that could dispose of this too large contribution. In our opinion, the suppression mechanism for both bulk and brane (SM and possibly BSM) contributions should be driven by a global renormalization group (RG) flow that takes into account the whole $(4+1)$D theory, including the SM in the 3-brane. The UV-complete theory (string theory/quantum gravity) should provide the boundary (UV) value of the vacuum energy (more generally the boundary values of all the physical parameters), and RG equations should dictate how the measured vacuum energy is reached in the IR [32]. Assessing the presence/absence of UV-sensitive terms in loop calculations is crucial to determine this UV-IR flow. Work is in progress in this direction [33].
Acknowledgments
We thank G. Dvali and M. Scalisi for useful discussions. The work of CB has been supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567), and by the European Union – Next Generation EU through the research grant number P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Università e Ricerca (MUR). The work of VB, FC and AP is carried out within the INFN project QGSKY.
I Appendix A - Dark Dimension scenario
For the reader’s convenience, in this Appendix we provide an expanded version of the introduction to the DD scenario given in the text. The proposal is formulated combining observational data and phenomenological bounds with swampland conjectures. It is based on the assumption that our universe lies in a unique corner of the quantum gravity landscape. This region is determined by the distance conjecture [34], according to which at large distances in the moduli field space $\phi$ the scale $\mu_{tow}$ of one of the towers present in these asymptotic regions becomes exponentially small, $\mu_{tow}\sim e^{-\alpha|\phi|}$, with $\alpha$ positive $\mathcal{O}(1)$ constant. When this conjecture is applied to (A)dS vacua [^9] The AdS version of the conjecture is supported by a larger amount of arguments in string theory. The relation (35) is nevertheless assumed to be valid also in dS space., the relation
$\mu_{tow}\sim\left|\frac{\Lambda_{cc}}{M_{P}^{4}}\right|^{\alpha}M_{P}$ (35)
is established [35].
Combining indications from one-loop string calculations with the Higuchi bound [4], the authors of [2] restrict $\alpha$ to the range
$\frac{1}{4}\leq\alpha\leq\frac{1}{2}\,.$ (36)
Moreover, according to the emergent string conjecture [36, 37], they observe that in these asymptotic regions only two possible kinds of towers are expected: either a tower of string excitations, or one of KK states.
Since their goal is to apply these conjectures/results to our universe, they begin by observing that a light tower of states causes significant deviations from the gravitational Newton’s force law at the energy scale  [^10] For a KK tower this is due to gravity propagating on the extra dimensions, while for the string case local physics breaks down at the string scale. $\mu_{tow}$, and that torsion balance experiments provide the strongest bounds to deviations of the $1/r^{2}$ law. Noticing that this law was verified down to scales around  30 ${\rm\mu m}$[38], which implies that $\mu_{tow}$ has to satisfy the bound
$\mu_{tow}\geq 6.6\,\,{\rm meV},$ (37)
and that this value of $\mu_{tow}$ is close to the energy scale associated to the cosmological constant, $\Lambda_{cc}^{1/4}\sim 2.31\,\,{\rm meV}$, they conclude that $\mu_{tow}$ must satisfy the relation
[…]
otherwise we would have already detected deviations from Newton’s law. They then observe that the only way to make the “experimental bound” (38) consistent with the “theoretical swampland bound” (35) is to take $\alpha=1/4$ (the quotation marks are to indicate that we are using their words). This leads them to conclude that: (i) “there is a tower of states starting at the neutrino scale in our universe”; (ii) “this tower is either a light string tower or a light KK tower”. Concerning this latter point, they observe that “since we can describe physics above the neutrino scale with Effective Field Theory” (our italics), the string scenario is “ruled out experimentally”, as no experiment has ever detected a string spectrum at these scales.
The authors conclude that experiments tell us that the light tower must be a KK tower, $\mu_{tow}=m_{{}_{\rm KK}}$, so that (in their words) “the only possibility left” is that of a “scenario of decompactification”, where (35) becomes $\Lambda_{cc}\sim m_{{}_{\rm KK}}^{4}$, that is nothing but (4). The last step in the construction of their proposal is the determination of the number of extra compact dimensions. To this end, they compare bounds on $m_{{}_{\rm KK}}$ that come from calculations of the heating of neutron stars due to the surrounding cloud of trapped KK gravitons [5, 6] with the value of $m_{{}_{\rm KK}}$ given in (4), that leads to the conclusion that there can be only one extra dimension of size $\sim\mu{\rm m}$.
Therefore, assuming that our universe lies in a unique corner of the quantum gravity landscape, and combining swampland conjectures with phenomenological and observational data, the Dark Dimension scenario predicts that at low energies physics is described by a $(4+1)$D EFT with a compact dimension of micrometer size.
II Appendix B - Standard Model cutoff
In this Appendix we derive the relation between the cutoff $\Lambda$ of a $(4+1)$D theory and the $4$D cutoff $\Lambda_{\rm SM}$ of the Standard Model localized on a 3-brane. Let us consider the $(4+1)$D theory (with compact space dimension in the shape of a circle of radius $R$) defined by
$\mathcal{S}=\mathcal{S}_{\rm grav}+\mathcal{S}_{\rm{mat}},$ (39)
where
[…]
is the $(4+1)$D Einstein-Hilbert action and as an example for the matter action we take
[…]
with $\hat{\Phi}$ a $(4+1)$D scalar field that obeys the boundary condition $\hat{\Phi}(x,z+2\pi R)=\hat{\Phi}(x,z)$. We indicate with $x$ the $4$D coordinates and with $z$ the coordinate along the compact dimension. Using the signature (+,,,,)(+,-,-,-,-)( + , - , - , - , - ), the $(4+1)$D metric is parametrized as
$\hat{g}_{{}_{MN}}=\begin{pmatrix}e^{2\alpha\phi}g_{\mu\nu}-e^{2\beta\phi}A_{\mu}A_{\nu}&e^{2\beta\phi}A_{\mu}\\ e^{2\beta\phi}A_{\nu}&-e^{2\beta\phi}\,,\end{pmatrix}$ (42)
where $A_{\mu}$ is the so called graviphoton and $\phi$ the radion field. Considering only zero modes for $\hat{g}_{{}_{MN}}$, i.e.  $g_{\mu\nu}(x)$, $A_{\mu}(x)$ and $\phi(x)$ only depend on $x$, and integrating over $z$, for the 4D gravitational action $\mathcal{S}^{(4)}_{\rm grav}$ we get [39]
$\mathcal{S}^{(4)}_{\rm grav}=\frac{1}{2\kappa^{2}}\int\mathrm{d}^{4}x\,\sqrt{-g}$
×$\times\left[\mathcal{R}-2e^{2\alpha\phi}\hat{\Lambda}_{cc}+{2}\alpha\Box\phi+\frac{(\partial\phi)^{2}}{2}-\frac{e^{-6\alpha\phi}}{4}F^{2}\right],$ (43)
where the $4$D constant $\kappa=M_{P}^{2}$ is related to the $(4+1)$D $\hat{\kappa}=\hat{M}_{P}^{3}$ through the relation $\kappa^{2}=\hat{\kappa}^{2}/(2\pi R)$. The fields $\phi$ and $A_{\mu}$ in the above equation are dimensionless (dimensionful fields are obtained through the redefinition $\phi\to\phi/(\sqrt{2}\kappa),\,A_{\mu}\to A_{\mu}/(\sqrt{2}\kappa)$), and we used $02\alpha+\beta=$ 0. The canonical kinetic term in (II) for the radion field is obtained taking $\alpha={1}/{\sqrt{12}}$.
Considering the Fourier decomposition of $\hat{\Phi}(x,z)$, for the $4D$ matter action (41) we have
$\mathcal{S}^{(4)}_{\rm mat}=\int d^{4}x\sqrt{-g}\,\,$ (44)
×$\times\sum_{n}\left[\left|D\varphi_{n}\right|^{2}-\left(e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}m^{2}+e^{\sqrt{6}\frac{\phi}{M_{P}}}\frac{n^{2}}{R^{2}}\right)\left|\varphi{{}_{n}}\right|^{2}\right],$
where $D_{\mu}\equiv\partial_{\mu}-i\left({n}/{R}\right)A_{\mu}$, and $\varphi_{n}(x)$ are the KK modes of $\hat{\Phi}(x,z)$.
Taking a constant background for the radion (that for notational simplicity we continue to indicate with $\phi$) and the trivial background for $A_{\mu}$, the metric (42) becomes
$\hat{g}^{0}_{{}_{MN}}=\begin{pmatrix}e^{\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}\eta_{\mu\nu}&0\\ 0&-e^{-2\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}\end{pmatrix}.$ (45)
From (44) we define the $\phi$-dependent radius $R_{\phi}\equiv R\,e^{-\sqrt{\frac{3}{2}}\frac{\phi}{M_{P}}}$. With such a definition, we immediately see that, when computing radiative corrections, the $(4+1)$D momentum $\hat{p}\equiv(p,n/R)$ is cut as
$\hat{p}^{2}=e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{P}}}\left(p^{2}+\frac{n^{2}}{R_{\phi}^{2}}\right)\leq\Lambda^{2}.$ (46)
This latter equation is conveniently rewritten as
$p^{2}+\frac{n^{2}}{R_{\phi}^{2}}\leq\Lambda_{\phi}^{2},$ (47)
where we defined $\Lambda_{\phi}\equiv\Lambda\,e^{\frac{1}{\sqrt{6}}\frac{\phi}{M_{P}}}$. In terms of the dimensionless $\phi$ of (42) and (II), and before using $\alpha=1/\sqrt{12}$, it is $\Lambda_{\phi}=e^{\alpha\phi}\Lambda=m_{{}_{\rm KK}}^{1/3}R^{1/3}\Lambda$.
As $p^{2}$ in (47) is nothing but the modulus of the four-momentum on the brane, this equation tells us that $\Lambda_{\phi}$ is the cutoff $\Lambda_{\rm SM}$ of the SM (or more generally of the BSM model that lives on the 3-brane, where fields have $n=0$). Therefore:
$\Lambda_{\rm SM}=\Lambda_{\phi}=\Lambda\,e^{\frac{1}{\sqrt{6}}\frac{\phi}{M_{P}}}.$ (48)
Finally, as the DD scenario is realized for negative values of $\phi$, from (48) we see that $\Lambda_{\rm SM}\leq\Lambda$, i.e.  the SM cutoff is lower than the cutoff of the $(4+1)$-dimensional EFT that implements the DD scenario.
Before closing this Appendix, let us note that here we considered a spherical cutoff. Naturally, we can make a different choice, taking for instance a cylindrical cutoff [27]
$p^{2}\leq\Lambda_{\phi}^{2}\qquad{\rm and}\qquad\frac{n^{2}}{R_{\phi}^{2}}\leq\Lambda_{\phi}^{2}.$
This choice, that is closer to what is typically done when using the species scale $\Lambda_{\rm sp}$ as the cutoff [26], does not change the above considerations.

### v2

@@
 In [1] we showed that the right hand side of (5) also contains UV-sensitive terms (see the first line in the right hand side of Eq.
-(Dark Dimension and the Effective Field Theory limit) below), that were missed in previous literature.
+(Dark Dimension and the Effective Field Theory limit) below), that were missed in previous literature [^3] The presence of similar UV-sensitive terms in the calculation of the effective potential of a scalar theory was first hinted in [24].
+Moreover, it was shown in [25] that the presence of Fayet-Iliopulos terms induces further divergences..
 Therefore, the matching between the prediction (4) for the vacuum energy $\Lambda_{cc}$, that comes from swampland conjectures in string theory, and the corresponding result in the low energy field theory limit is not as straightforward as it was thought on the basis of the result (5).
@@
 However, we think that such a delicate issue cannot be trivially shortcut with the choice $\Lambda=m_{{}_{\rm KK}}$ considered in [3] and discussed above.
-Finally, we would like to spend one word on the general framework in which in our opinion the DD scenario should be framed.
+Finally, we would like to spend one word on the general framework where, in our opinion, the DD scenario should be framed.
 In [1] and in the present work we considered the contribution to the vacuum energy that comes from the bulk.
-The reason is that in the paper that contains the DD proposal [2] as well as in the paper [3] that attempts to confute the results of our work [1] this is the only part of the vacuum energy taken into account.
-Naturally, there is also the contribution coming from the Standard Model, that lives on a 3-brane.
-Since the SM physical cutoff $\Lambda_{\rm SM}$ is at least of the TeV order, the latter cannot be ignored and viability of the DD scenario is subject to the existence of a physical mechanism that could dispose of this too large contribution.
-In our opinion, the suppression mechanism for both bulk and brane (SM and possibly BSM) contributions should be driven by a global renormalization group (RG) flow that takes into account the whole $(4+1)$D theory, including the SM in the 3-brane.
-The UV-complete theory (string theory/quantum gravity) should provide the boundary (UV) value of the vacuum energy (more generally the boundary values of all the physical parameters), and RG equations should dictate how the measured vacuum energy is reached in the IR [32].
-Assessing the presence/absence of UV-sensitive terms in loop calculations is crucial to determine this UV-IR flow.
+The reason is that in the paper that contains the DD proposal [2], as well as in the paper [3] that attempts to confute the results of our work [1], this is the only part of the vacuum energy taken into account.
+Naturally, there is also a contribution from the Standard Model, that, as already said, lives on a 3-brane.
+Since the SM physical cutoff $\Lambda_{\rm SM}$ is at least of the TeV order, this contribution cannot be ignored.
+In this respect, it is important to stress that it has been already objected in the past (well before the formulation of the DD proposal) that the SM contribution to the vacuum energy seems to invalidate any attempt to get a small cosmological constant in theories with compact extra dimensions [34, 35].
+The viability of the DD scenario is certainly subject to the existence of a physical mechanism that could dispose of this too large contribution.
+We expect that, for both bulk and brane contributions, such a mechanism (if any) should be provided by the piling up of quantum fluctuations in the UV $\to$ IR renormalization group (RG) flow that takes into account the whole $(4+1)$D theory, including the SM in the 3-brane.
+In such a framework, the UV-complete theory (string theory/quantum gravity) should provide the boundary (i.e.
+UV) value of the vacuum energy (more generally the boundary values of all the physical parameters), so that RG equations should dictate how the measured vacuum energy is reached in the IR (see [36] for examples where this mechanism is realized in $d=3$ and $d=4$ scalar theories).
 Work is in progress in this direction [37].
 Acknowledgments
-We thank G.
-Dvali and M.
+We thank S.
+Abel, G.
+Dvali, H.
+P.
+Nilles and M.
 Scalisi for useful discussions.
-The work of CB has been supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567), and by the European Union – Next Generation EU through the research grant number P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Università e Ricerca (MUR).
+The work of CB has been partly supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567), and partly by the European Union – Next Generation EU through the research grant number P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Università e Ricerca (MUR).
 The work of VB, FC and AP is carried out within the INFN project QGSKY.